Carl is 7 years older than Anne. Fifteen years from now Carl will be 33 years less than twice Anne's age at that time. How old is Carl now?

Answer by addingup(3677)

C = A+7 We'll use this value for C next
and
C+15 = 2(A+15)-33 Substitute for C on the left:
A+7+15 = 2A+30-33
A+22 = 2A-3 Subtract A on both sides and add 3 on both sides:
A-A+22+3 = 2A-A-3+3 Simplify:
25 = A
or, more elegantly:
A = 25 Anne is 25 and so Carl is 25+7 = 32
-------------------------------------------
In 15 years:
Carl: 32+15 = 47
Anne:25+15 = 40
At that time, 15 years from now, Carl will be 2A-33. Let's see:
47 = 2(40)-33
47 = 80-33 Correct
:
John